Rose's garden is in the shape of a trapezoid. if the height of the trapezoid is 16m, one base is 20m, and the area is 224m squared, what is the length of the other base?

Answer by stanbon(75887)

Formula: Area = (1/2)(height)(b1+b2) where b1 and b2 are the parallel bases.
(b1+b2) = 2(base)/(height)
(b+20) =2(224)/16
b+20 = 28
b = 8cm